OpenStudy (anonymous):

ALG 2 Quadratic Formula / Completing the Square What are the solutions of 2x^2 – 6x = –8? a) -4 ± i√26 ÷ 8 b) -1 ± i√5 ÷ 2 c) -1 ± 2i√2 d) 3 ± i√7 ÷ 6

OpenStudy (anonymous):

none of the above

OpenStudy (anonymous):

they are \[\frac{3\pm\sqrt{7}i}{2}\]

OpenStudy (anonymous):

that's what i thought but it is multiple choice and my answer was 3 ± i√7 ÷ 2

OpenStudy (anonymous):

your answer is right. it matches none of the above

OpenStudy (anonymous):

Thanks, so I'm just gonna go with d, maybe they messed up putting the answer choices in...

OpenStudy (amistre64):

2x^2 – 6x = –8? ; /2 x^2 -3x = -4 ; complete square (x-3/2)^2 = -16+9/4 x = 3/2 +- sqrt(-7)/2 ..... maybe